OpenStudy (anonymous):

3/4x-6=9 Solve the equation

OpenStudy (anonymous):

its 3 over 4 not divided

OpenStudy (anonymous):

\[\frac{3}{4}x-6=9\] what happens if you add 6 to both sides?

OpenStudy (anonymous):

ohhhh you get 15 and the other -6 cancles out

OpenStudy (anonymous):

good job and you're left with \[\frac{3}{4}x=15\]

OpenStudy (anonymous):

now you want to get rid of the 3/4 what can you multiply by to get rid of it

OpenStudy (anonymous):

right so you divide next? idk im confused

OpenStudy (anonymous):

the 4?

OpenStudy (anonymous):

ok what happens when you multiply by 4 on both sides?

OpenStudy (anonymous):

you get 60.

OpenStudy (anonymous):

and the other side is?

OpenStudy (anonymous):

16?

OpenStudy (anonymous):

??? \[\frac{3}{4}*4 = ???\] cancel out the 4 and you're left with?

OpenStudy (anonymous):

ohhh sorry then your left with 3x?

OpenStudy (anonymous):

exactly now you have 3x=60 you want to get rid of the 3 what do you do now?

OpenStudy (anonymous):

x=20! :)

OpenStudy (anonymous):

yay, you actually picked up on this quite quickly ok just remember, if you ever do anything (add subtract multiply divide) on one side, you must do it on the other side otherwise it wouldnt be a balanced equation anymore

OpenStudy (anonymous):

okay can you help me with two more?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Thanks! :) okay ummm -2/3y+12=-8 again -2 over 3

OpenStudy (anonymous):

\[-\frac{2}{3}y+12 = -8\] ok what do you think you should do first?

OpenStudy (anonymous):

ummm subtract the 12 on both sides?

OpenStudy (anonymous):

good and you will get \[-\frac{2}{3}y = -20\] ok what next?

OpenStudy (anonymous):

umm multiply the 3?

OpenStudy (anonymous):

it doesnt matter if you multiply by 3 or divide by 2 first but yes if you multiply by 3, then what will you get?

OpenStudy (anonymous):

Wait its a negative 3?

OpenStudy (anonymous):

if its not then the answer will be negative 60?

OpenStudy (anonymous):

sorry if you're still here multiply by -3 and you get 2y=60 then y=30

OpenStudy (anonymous):

os crashed on me